OpenStudy (anonymous):

{Will give medal} Quick question: How do I finish this problem? A rectangular field will be fenced on all four sides. Fencing for the north and south sides costs $10 per foot and fencing for the other two sides costs $9 per foot. What is the maximum area that can be enclosed for $6000? 20L+18W=6000 18W = 6000 - 20L W = 1000/3 - 10L/9 A = L*W A = L (1000/3 - 10L/9) A = 1000L/3 - 10L^2/9

OpenStudy (anonymous):

Take the derivative of \(A\) with respect to \(L\): \[\frac{dA}{dL}=\frac{1000}{3}-\frac{20L}{9}\] and find the critical points (i.e. solve \(\dfrac{dA}{dL}=0\) for \(L\)).